# Simulation of Hypergraph Algorithms with Looped Transformers

### v1

@@ 5.1 Iteration of Visiting Hyperedges
Let the operation of visiting hyperedges iteratively be defined as described in Algorithm 3. This operation requires 1 increment operation, which requires single layer transformer to construct the following from Lemma 4.2, 2 compare operations, which require 3 layers transformer to construct the following from Lemma 4.4, 3 selection operations, which require 3 layers transformer to construct following from Lemma 4.1, 1 AND operation which require single layer transformer to construct following from Lemma 4.8, 1 read-from-incident-matrix operation which requires single layer transformer to construct following from Lemma 4.7, 1 write-scalar-to-column operation which requires single layer transformer to construct following from Lemma 4.5, and 1 trigger-termination operation which require single layer transformer to construct following from Lemma 4.6. The whole algorithm can be constructed by
@@
5.2 Iteration of Minimum Value
Lemma 5.2 (Get minimum value).
If the following conditions hold:
  • Let the transformer be defined as Definition 3.11.
@@ 5.2 Iteration of Minimum Value
Following from Lemma 4.1, Lemma 4.2, Lemma 4.3, Lemma 4.4, Lemma 4.5, and Lemma 4.6, we can perform all the operation of Algorithm 2, thus we finish the proof. ∎
5.3 Dijkstra’s Algorithm
Theorem 5.3 (Dijkstra’s Algorithm on hypergraph).
@@ 5.3 Dijkstra’s Algorithm
Let Dijkstra’s algorithm be considered as described in Algorithm 4. Following from Lemma 5.1 and Lemma 5.2, the operation of iteratively visiting vertices and hyperedges requires 18 layers, as established in these lemmas. For the remaining part of the algorithm, the operations include 4 selection operations, 1 add operation, 1 compare operation, 1 repeat AND operation, 1 write scalar to column operation, and 1 trigger termination operation. According to Lemma 4.1, Lemma 4.10, Lemma 4.9, Lemma 4.4, Lemma 4.5, and Lemma 4.6, these operations together require 9 layers. Therefore, the total number of layers required to simulate Dijkstra’s algorithm is:
$18+9=27$
layers of the transformer. ∎
6 Main Results
6.1 Degradation
We observe that [13] presents the running results of several algorithms on graphs, which primarily include Dijkstra’s algorithm for the shortest path, Breadth-First Search (BFS), and Depth-First Search (DFS). [13] provides a general representation of such algorithms in Algorithm 7 of their paper. Given that Dijkstra’s algorithm for the shortest path, BFS, and DFS can be straightforwardly extended to hypergraphs [21] by identifying the shortest hyperedge between two nodes and treating it as the distance between those nodes, we propose a degradation mechanism in Algorithm 3. This mechanism allows dynamic access to the shortest hyperedge between two vertices while maintaining the feature dimension of the Looped Transformer at $O(1)$. Using this mechanism, we can extend Dijkstra’s algorithm for the shortest path, BFS, and DFS from graphs to hypergraphs. We formally state the theorem regarding the degradation mechanism as follows:
Theorem 6.1 (Degradation).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.11, consisting of 10 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer simulates the degradation operation, i.e., the degradation of the incident matrix to the adjacency matrix while retaining only the shortest hyperedge between two vertices, for weighted hypergraphs, supporting up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges.
Proof.
Using the same construction as Lemma 5.1, we finish the proof. ∎
6.2 Helly
The Helly property in a hypergraph refers to a specific condition in the family of its hyperedges. A hypergraph is said to satisfy the Helly property if, for every collection of its hyperedges, whenever the intersection of every pair of hyperedges in the collection is non-empty, there exists at least one hyperedge in the collection that intersects all the others. In other words, if the pairwise intersections of a set of hyperedges are non-empty, there must be a single hyperedge that intersects every hyperedge in the set. This property is named after Helly’s theorem in convex geometry, which inspired its application in combinatorial settings like hypergraphs.
We have reformulated the algorithm for determining whether a hypergraph possesses the Helly property, as presented in Algorithm 3 of [5], into a form that the Looped Transformer can execute. This reformulation is represented by our Algorithm 5. We now state the following theorem:
Theorem 6.2 (Helly).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.11, consisting of 11 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer simulates the Helly algorithm (Algorithm 5) for weighted hypergraphs, handling up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges.
Proof.
Let the degradation operation be defined as described in Algorithm 3. This operation requires 3 increment operations, 5 selection operations, 1 AND operation, 1 compare operation, and 1 read-from-incident-matrix operation. By applying Lemma 4.2, Lemma 4.1, Lemma 4.8, Lemma 4.4, and Lemma 4.7, each of these operations can be constructed using a single-layer transformer, as detailed in the corresponding lemmas. The total number of layers required for the entire degradation operation is:
$3+5+1+1+1=11$
Thus, the degradation operation can be constructed using 11 layers of transformer. This completes the proof. ∎
7 Conclusion

### v2

@@ Appendix D Missing Proof in Simulation
Let the operation of visiting hyperedges iteratively be defined as described in Algorithm 2. This operation requires 1 increment operation, which requires single layer transformer to construct the following from Lemma C.2, 2 compare operations, which require 3 layers transformer to construct the following from Lemma C.3, 3 selection operations, which require 3 layers transformer to construct following from Lemma C.1, 1 AND operation which require single layer transformer to construct following from Lemma C.8, 1 read-from-incident-matrix operation which requires single layer transformer to construct following from Lemma C.7, 1 write-scalar-to-column operation which requires single layer transformer to construct following from Lemma C.5, and 1 trigger-termination operation which require single layer transformer to construct following from Lemma C.6. The whole algorithm can be constructed by
@@
Appendix E Missing Proof in Main Results
E.1 Degradation
Theorem E.1 (Degradation, formal version of Theorem 4.1).
A looped transformer $h_{T}$ defined in Definition 3.11 exists, consisting of 10 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer can simulate the degradation operation (Algorithm 2) for hypergraphs, supporting up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges, where $\widehat{\delta}$ defined in Definition 3.12 is the nearest representable approximation of the minimum increment angle.
@@ E.1 Degradation
Using the same construction as Lemma 6.1, we can show that 10 layers can iterate the algorithm. Furthermore, for an incident matrix, we require its dimensions to be within the smallest computable precision. Therefore, it is necessary to ensure that there are at most $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges. ∎
E.2 Helly
Theorem E.2 (Helly, formal version of Theorem 4.2).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.10, consisting of 11 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer can simulate the Helly algorithm (Algorithm 5) for weighted hypergraphs, handling up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges, where $\widehat{\delta}$ defined in Definition 3.12 is the nearest representable approximation of the minimum increment angle.
@@ E.2 Helly
This algorithm requires 3 increment operations, 5 selection operations, 1 AND operation, 1 compare operation, and 1 read-from-incident-matrix operation. By applying Lemma C.2, Lemma C.1, Lemma C.8, Lemma C.3, and Lemma C.7, each of these operations can be constructed using a single-layer transformer, as detailed in the corresponding lemmas. The total number of layers required for the entire degradation operation is: $3+5+1+1+1=11$. Thus, the degradation operation can be constructed using 11 layers of transformer. Furthermore, for an incident matrix, we require its dimensions to be within the smallest computable precision. Therefore, it is necessary to ensure that there are at most $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges. ∎